Sami spent 1/3 of his money on books and the remaining money on 5 identical presents. What fraction of the money did he spend on each of the presents?
Answer by Theo(13342)

sami spent 1/3 of his money on books and the remaining money on 5 identical presents.

what fraction of his money did he spend on each of the presents?

since sami spent 1/3 of his money on books, then his remaining money must have been 2/3 of his money because 1/3 + 2/3 = 1.

2/3 of his money divided by 5 is equal to 2/15 of his money.

that's how much of his money he spent on each of the 5 identical presents.

5 * 2/15 + 1/3 is equal to 10/15 + 1/3.
1/3 * 5/5 = 5/15.
10/15 + 1/3 is equal to 10/15 + 5/15 which is equal to 15/15 which is equal to 1.

solution looks good.

he spent 2/15 of his money on each of the presents.